# Applying a STEM Ways of Thinking Framework for Student-generated Engineering Design-based Physics Problems

### v1

@@ 4.4 Data Analysis
Student-groups recorded their discussions in the laboratory or corridors, resulting in background noise that made it extremely challenging to uniquely associate conversation segments with individual speakers. Additionally, the final report in week 14 was a collective group output. These factors, along with the fact that students worked in groups in all labs, influenced our decision to choose the “grain size” or “unit of analysis” [10, p.10] to be that of a group. The audio recordings, ranging in duration from 5 to 9 minutes with an average of about 6 minutes, were all manually transcribed by the lead author. The transcript lengths ranged from about 600 to 1200 words each, with the average being about 850.
@@
Throughout the process, the lead author took detailed handwritten notes, reviewing the data multiple times and documenting reflections and coding decisions. In line with his previous studies [9, 79, 7, 5], the lead author, as his his wont, adopted a traditional pen and paper approach coupled with Word and Excel as his “codus operandi” [92, p.26], which facilitated direct engagement with the data, allowing for valuable insights to emerge through close interaction with the material.
Data (transcripts and written reports) from all 14 teams is used for each research question. However, as will be explained in the Findings and Discussion section, the portion of the data selected for each research question varied based on the nature of the question and the suitability of the data. Details are provided in Table 4.
Table 4: Data from all 14 teams were used for each research question. However, the specific portion of data selected for each question varied depending on the nature of the question and the suitability of the data. (RQ – Research Question; PF – Problem Framing; DBT – Design-based Thinking; SBT – Science-based Thinking; MBT – Math-based Thinking; MER – Metacognitive Reflection; CT – Computational Thinking)
RQ Aspect Portion of data selected Justification provided in
@@
Building on his prior studies [5, 4], the lead author, with assistance from another researcher, coded the transcripts based on five Ways of Thinking in our WoTEDP framework (see Table 5). Following our prior approach [9], we used flexible coding units, considering student discussion complexity and group conversation styles [93, 10]. This choice was guided by communication patterns, thematic shifts, speaker transitions, contextual completeness, and our research questions [9]. As Elliot notes [94, p.2856]“there is no simple answer” to the choice of “the chunk of data” to be coded, and “it depends on the study and your aims within it”. Moreover, fixed text lengths might fail to capture the richness and diversity of these discussions, given that some groups engaged in dynamic exchanges in contrast to others that spoke more linearly. It is also worth noting that multiple codes (code co-occurrence) [95][94, p.2853] could be applied to a single segment. To account for the varying transcript lengths and unequal textual segments, we calculated fractional coding frequencies for each team and averaged them across all groups. Given the qualitative nature of the study and our focus on broader trends rather than statistical significance of differences in code percentages, error bars have not been included [9] in Figure 3.
@@
Table 5: The table below shows the Ways of Thinking codes for transcripts, code descriptions, example quotes, and code justification. (DBT–Design Based Thinking; SBT–Science Based Thinking; MBT–Math Based thinking; MER–Metacognitive Reflection; CT - Computational Thinking.)
@@
As to why we describe this stage in detail despite not relying on it extensively, the reason is methodological: it was a crucial first step in our analysis. This stage provided important insights that motivated the development of a more layered coding approach. Moreover, this initial method proves inadequate for analyzing written reports, where the data is more complex, integrating text, Python code, figures, graphs, equations, and tables. Furthermore, the diversity in teams’ writing (inclusive of drawing and coding) styles added yet another layer to the already intricate and multifaceted nature of the data, which O’Dwyer terms “a messy but attractive nuisance” [100]. As in our previous study [9], we conducted a second round of coding, which we describe in the following section.
@@
the lead author transitioned to a “constructive (during the process)” [103, p.15] procedure as opposed to a traditional “evaluative (post hoc)” [103, p.15] inter rater reliability (IRR) test. Taking guidance from Cascio et al. [104], we developed a team-based approach which we term Team-Based Efforts to Enhance Reliability (TBEER). Over a two-month period, two coauthors assisted in the analysis, meeting with the lead author at least once a week for an hour. The lead author coded all 14 transcripts and 14 reports, with one coauthor coding 7 transcripts and 7 reports, while the remaining transcripts and reports were coded by the other coauthor. Each team member independently coded the transcripts and written reports. They then compared, discussed, and debated their codes to iteratively reach a consensus. Cascio et al. term this an inter-coder consensus as opposed to inter coder reliability [104], while Linneberg and Korsgaard [105, p.268] term it inter-coder agreement. Through regular “peer debriefing” [103, p.14][87] sessions, the team collaboratively moved from data to findings. The lead author meticulously maintained an audit trail of the team’s deliberations through the study “to document the course of development of the completed analysis” [103, p.19].
@@
Given that our data is rooted in a physics laboratory incorporating ED-based tasks, our analysis and interpretation were deeply informed by our disciplinary understanding in Newtonian Mechanics. Although students occasionally referenced concepts from other areas of physics, their science-based thinking was primarily within the conceptual structure of Newtonian Mechanics [9].
@@
Adhering to Wa-Mbaleka’s guidance that qualitative research “may be messy, [but] we need to explain what we did, how we did it, and why we did it”[123, p.40], we strove to enhance transparency [116] and trustworthiness [124, 125] of our findings. To this end we share our coding sheets as space constraints would permit. Each subsection in our analysis includes a comprehensive explanation of our analytical approach, accompanied by the coding sheets we used. This is in line with the perspective of Devetak et al., who emphasize that the strength of good qualitative research lies in offering “sufficient details for the reader to grasp the idiosyncrasies of the situation” [15, p.83].
The reader may find it useful to refer our earlier studies [9, 6] in which we have addressed some of the methodological aspects of qualitative data analysis. Researchers interested in additional details on our analytical approach are welcome to contact us.
@@ 5 Findings & Discussion
One of the implicit goals of this study is to uncover patterns that can inform future pedagogical interventions. While we remain mindful of the limits of generalization, we aim to identify insights that hold local transferability within our educational setting [126, 127].
In this context, we find it important to follow the guidance offered by Guest et al. in describing the prevalence of a code among the 14 teams:
@@
As a rule of thumb, we use terms such as ‘a majority of the teams’ to refer to more than 8 teams, ‘about half’ for 6–8 teams, and a ‘smaller fraction’ for fewer than 6 teams. In our view, precise numbers may not be particularly relevant in our qualitative study, though the coding sheets do reveal them.
We emphasize that Ways of Thinking are not isolated elements. While they are interconnected, in this study, we focus on discussing these elements separately, though we occasionally highlight their interconnections. However, we will explore this aspect in depth in the third part of this series, focusing on Integrative Thinking. Readers may find it useful to refer to our recent study, where we examined design-science connections [9] in students’ thinking as they navigated an instructor-assigned ED task.
@@ 5.1 ED-problem Framing
In this section, before analyzing students’ solutions, we first examine aspects of their problem framing, given that the problem itself was framed by the students. Given the significance of problem framing, we provide an introductory discussion in Section (i). Section (ii) details our approach to examining the data, Section (iii) outlines key findings, and Section (iv) offers recommendations to support students in refining their problem framing.
@@
While giving students the agency to frame and solve their own problems is valuable, it comes with challenges. Most students are accustomed to solving well-defined textbook problems and rarely have opportunities to frame their own problems. To address this lacuna, we began the semester with an engineering design (ED) problem [9] that we had framed (see Table 2). Later in the semester, we introduced the student-generated ED problem task (see Tables 2 and 3). In week 10, student teams were asked to:
@@
Students were guided to progress towards their solutions from weeks 10 - 14. Students were encouraged to engage with TAs for guidance, while TAs were instructed to remain flexible and supportive, even when students set ambitious goals.
Our approach of guiding students toward framing a problem statement as a key competency aligns with the perspective that “developing competencies to solve ill-structured problems in teams is essential for engineering students” [130, p.460].
@@
As shown in Figure 4, we categorized the problem statement under design-based thinking. This classification aligns with the perspective that design thinking is a problem-solving approach that integrates creative thinking to drive innovation. A fundamental aspect of design thinking is problem framing, which shapes the direction of the design process [131, p.496].
@@ 11. Fungee Jumping:
Students began working on the Jupyter Notebooks for the ED problem in week 10, but our analysis focused solely on the final week’s PDF report. We relied on these reports rather than group discussions, as most teams prioritized solutions over revisiting their problem statements (see Table 4). This analysis examines the final product of problem framing, not its evolution over time. For coding results, refer to Appendix A. Figures 5 and 6 present the problem statements from two teams.
@@
Figure 5: Example problem statement from Team-05. The team identified stakeholders (engineering firm, companies), provided the reasoning behind their problem choice, outlined criteria (range and target location) and constraints (specific angle of elevation), modeled the spring to obey Hook’s law, provided metrics (numerical factors relevant to their problem), stated assumptions (negligible air resistance), and even had a name for their company. See Appendix A.
@@
Figure 6: Example problem statement from Team-09. The team identified the stakeholders (the company, users (children), parents), provided a reasoning behind their problem choice, stated criteria (material choice, safety) and constraints (safety), did not mention any numerical factors, and had a name for their company. See Appendix A.
@@
An interesting trend emerged: six teams assigned creative branding or marketing names to their designs, reflecting an entrepreneurial mindset [133] and engagement beyond mere technical problem-solving. Examples include Fungee Jumping (Team-11’s bungee jumping design), Hookean Bow (Team-05’s bow and arrow), and Duncan’s Toy Chest (Team-09’s bean bag toy). These names showcased students’ creativity and marketing instincts, highlighting an aspect of design thinking that, while valuable, may not be typically prioritized in a physics classroom. See Table 6.
@@
This example also highlights the variety of ways students chose to frame their problem statement.
About half the teams explicitly referenced assumptions and approximations in their problem statements. This may reflect an awareness of design complexities or an implicit plan to address these aspects later. Another challenge was distinguishing between assumptions and approximations, as students often used these terms interchangeably. For instance, Team-05 stated:
@@
In engineering design-based activities, the role of iterative thinking may need to be emphasized. Following a systematic approach to brainstorming [134] which we had outlined in our recent paper [9], may aid in guiding students to frame their problem statement. Valkenburg and Dorst [135, p.254] emphasize that reflective practices can help rethink and reframe problems, leading to better solutions.
It is also important for students to select problems that not only align with their personal interests but are simple enough so that meaningful progress is possible given limited time and resources. Actively seeking input from TAs—can enhance problem framing and engagement.
@@ 5.2 Design-Based Thinking (DBT)
While coding, we considered whether the team intended to refer to safety as a criterion rather than a constraint. One perspective is that safety is a constraint because it represents a non-negotiable requirement that the design must meet. We had a productive discussion on this issue during coding and leave it to the readers to continue the discussion. In our experience, this is not an isolated instance of the blurry difference between what may be called a criterion or a constraint. Despite this challenge, these are useful notions to adopt in a design process.
@@
Almost all teams cited sources or referenced online materials (see Figure 19 in Appendix B). A few teams mentioned gaining insights through discussions with their TA, while one team drew inspiration from classroom lectures. Additionally, some teams referenced national regulatory authorities, such as the National Institute of Standards and Technology (Team-08) and the National Highway Traffic Safety Administration (Team-02). While it may be argued that referencing is typically associated with academic rigor rather than design thinking, the practice of referencing plays a complementary role in guiding students towards justifying design choices, and informing iterative improvements. Although we did not find literature specifically addressing referencing within design thinking, we believe it serves as a valuable tool for refining ideas and linking them to established knowledge, thus promoting iterative design.
@@
Team-09 chose to include the numerical details and their python code, while Team-12 found it convenient to employ a python code (see Figure 7).
@@
We observed variations in how teams approached iterations. While some engaged in cursory or token iterations, a smaller fraction pursued meaningful refinements. Given that Week 14 was the final week of the semester, some teams may have experienced end-of-semester fatigue. Moreover, more iterations did not always lead to better solutions, highlighting the need for a more systematic approach to guiding students in the iteration process.
@@
To enhance student output in future iterations, we propose a more detailed rubric, though we acknowledge the risk of it becoming a checklist for students to simply tick off.

### v2

@@ 4.4 Data Analysis
Student-groups recorded their discussions in the laboratory or corridors, resulting in background noise that made it extremely challenging to uniquely associate conversation segments with individual speakers. Additionally, the final report in week 14 was a collective group assignment. These factors, along with the fact that students worked in groups in all labs, influenced our decision to choose the “grain size” or “unit of analysis” [10, p.10] to be that of a group. The audio recordings, ranging in duration from 5 to 9 minutes with an average of about 6 minutes, were all manually transcribed by the lead author. The transcript lengths ranged from about 600 to 1200 words each, with the average being about 850.
@@
Throughout the process, the lead author took detailed handwritten notes, reviewing the data multiple times and documenting reflections and coding decisions. In line with his previous studies [9, 79, 7, 5], the lead author, as is his wont, adopted a traditional pen and paper approach coupled with Word and Excel as his “codus operandi” [92, p.26], which facilitated direct engagement with the data, allowing for valuable insights to emerge through close interaction with the material.
Data (transcripts and written reports) from all 14 teams is used for each research question. However, as will be explained in the Findings and Discussion section, the portion of the data selected for each research question varied based on the nature of the question and the suitability of the data. Details are provided in Table 4. The data analysis evolved through two stages.
Table 4: Data from all 14 teams were used for each research question. However, the specific portion of data selected for each question varied depending on the nature of the question and the suitability of the data. (RQ – Research Question; PF – Problem Framing; DBT – Design-based Thinking; SBT – Science-based Thinking; MBT – Mathematics-based Thinking; MER – Metacognitive Reflection; CT – Computational Thinking)
RQ Aspect Portion of data selected for analysis Justification provided in
@@
Building on his prior studies [5, 4], the lead author, with assistance from another researcher, coded the transcripts based on five Ways of Thinking in our WoTEDP framework (see Table 5). Following our prior approach [9], we used flexible coding units, considering student discussion complexity and group conversation styles [93, 10]. This choice was guided by communication patterns, thematic shifts, speaker transitions, contextual completeness, and our research questions [9]. As Elliot notes [94, p.2856]“there is no simple answer” to the choice of “the chunk of data” to be coded, and “it depends on the study and your aims within it”. Moreover, fixed text lengths might fail to capture the richness and diversity of these discussions, given that some teams engaged in dynamic exchanges in contrast to others that spoke more linearly. It is also worth noting that multiple codes (code co-occurrence) [95][94, p.2853] could be applied to a single segment. To account for the varying transcript lengths and unequal textual segments, we calculated fractional coding frequencies for each team and averaged them across all teams. Given the qualitative nature of the study and our focus on broader trends rather than statistical significance of differences in code percentages, error bars have not been included [9] in Figure 3.
@@
Table 5: The table below shows the Ways of Thinking codes for transcripts, code descriptions, example quotes, and code justification. (DBT–Design-Based Thinking; SBT–Science-Based Thinking; MBT–Mathematics-Based thinking; MER–Metacognitive Reflection; CT - Computational Thinking.)
@@
The reason we describe this stage in detail despite not relying on it extensively is methodological. This stage – a crucial first step in our analysis – provided important insights that motivated the development of a more layered coding approach. Moreover, this initial method proves inadequate for analyzing written reports, where the data is more complex, integrating text, Python code, figures, graphs, equations, and tables. Furthermore, the diversity in teams’ writing (including diagrams and coding) styles added yet another layer to the already intricate and multifaceted nature of the data, which O’Dwyer terms “a messy but attractive nuisance” [100]. As in our previous study [9], we conducted a second round of coding, which we describe in the following section.
@@
the lead author transitioned to a “constructive (during the process)” [103, p.15] procedure as opposed to a traditional “evaluative (post hoc)” [103, p.15] inter rater reliability (IRR) test. Taking guidance from Cascio et al. [104], we developed a team-based approach which we term Team-Based Efforts to Enhance Reliability (TBEER). Over a two-month period, two coauthors (Nikhil Borse [NB] and Winter Allen [WA]) assisted in the analysis, meeting with the lead author at least once a week for about an hour. While lead author coded all 14 transcripts and 14 reports, NB coded 7 transcripts and 7 reports, and the remaining transcripts and reports were coded by WA. The lead author, NB, and WA independently coded the transcripts and written reports. They then compared, discussed, and debated their codes to iteratively reach a consensus. Cascio et al. term this an inter-coder consensus as opposed to inter coder reliability [104], while Linneberg and Korsgaard [105, p.268] term it inter-coder agreement. Through regular “peer debriefing” [103, p.14][87] sessions, the team collaboratively moved from data to findings. The lead author meticulously maintained an audit trail of the team’s deliberations through the study “to document the course of development of the completed analysis” [103, p.19].
@@
Given that our data is rooted in a physics laboratory incorporating ED-based tasks, our analysis and interpretation were deeply informed by our disciplinary understanding of Newtonian Mechanics. Although students occasionally referenced concepts from other areas of physics, their science-based thinking was primarily within the conceptual structure of Newtonian Mechanics [9].
@@
Adhering to Wa-Mbaleka’s guidance that qualitative research “may be messy, [but] we need to explain what we did, how we did it, and why we did it”[123, p.40], we strove to enhance transparency [116] and trustworthiness [124, 125] of our findings. To enhance transparency, we integrate our methodology throughout the manuscript. While broad details are introduced is this section, for clearer communication, finer aspects are presented in the next section on Findings & Discussions, which is directly related to the research questions.
Additionally, some aspects of methodology are revisited in the reflexivity statement in Section 7. The reader may also find it useful to refer to our earlier studies [9, 6] or contact us directly.
@@ 5 Findings & Discussion
In our pursuit of methodological transparency, specific to each research question, we amplify the analytical details in Section (ii), Section (i), Section (i), Section (i), Section (i), and Section (i). Although this might appear to be an overlap with the Methods section, this is in line with the recommendation of Devetak et al. that the strength of good qualitative research lies in offering “sufficient details for the reader to grasp the idiosyncrasies of the situation” [15, p.83].
To further orient the reader, we highlight a key consideration in addressing the research questions: uncovering patterns that may inform future pedagogical interventions. While we remain mindful of the limits of generalization, we aim to identify insights that hold local transferability within our educational setting [126, 127]. To this end, each subsection concludes with a ‘Moving Forward’ note, offering suggestions for improvement in our next iteration of pedagogical intervention.
In presenting our findings, we maintain the qualitative focus by following the guidance of Guest et al. in describing the prevalence of a code among the 14 teams:
@@
As a rule of thumb, we use terms such as ‘a majority’ or ‘a significant number’ or ‘most’ to refer to more than 8 teams, ‘about half’ for 6–8 teams, and a ‘smaller fraction’ or ‘only a few’ for fewer than 6 teams. In our view, precise numbers may not be particularly relevant in our qualitative study, though the coding sheets (see Appendix AAppendix B, and Appendix C) do reveal them.
We reiterate that Ways of Thinking are not isolated elements. While they are interconnected, in this study, we focus on discussing these elements separately, though we occasionally highlight their interconnections. However, we will explore this aspect in depth in the third part of this series, focusing on Integrative Thinking. Readers may find it useful to refer to our recent study, where we examined design-science connections [9] in students’ thinking as they navigated an instructor-assigned ED task.
@@ 5.1 ED-problem Framing
In this section, before analyzing students’ solutions, we first examine aspects of their problem framing, given that the problem itself was framed by the students. To emphasize the significance of problem framing, we provide an introductory discussion in Section (i). Section (ii) details our approach to examining the data, Section (iii) outlines key findings, and Section (iv) offers recommendations to support students in refining their problem framing.
@@
While giving students the agency to frame and solve their own problems is valuable, it comes with challenges. Most students are accustomed to solving well-defined textbook problems and rarely have opportunities to frame their own problems [130, 131]. To address this lacuna, we began the semester with an engineering design (ED) problem [9] that we had framed (see Table 2). Later in the semester, we introduced the student-generated ED problem task (see Tables 2 and 3). In week 10, student teams were asked to:
@@
Students were guided to progress towards their solutions from weeks 10 - 14. Students were encouraged to engage with TAs for guidance, while TAs were instructed to remain flexible and supportive, even when students set ambitious goals (see Section 2).
Our approach of guiding students toward framing a problem statement as a key competency aligns with the perspective that “developing competencies to solve ill-structured problems in teams is essential for engineering students” [132, p.460].
@@
As shown in Figure 4, we categorized the problem statement under design-based thinking. This classification aligns with the perspective that design thinking is a problem-solving approach that integrates creative thinking to drive innovation. A fundamental aspect of design thinking is problem framing, which shapes the direction of the design process [133, p.496]. It may be noted that since students are expected to apply physics concepts, an overlap with science-based thinking is not only inevitable, but indeed welcome.
@@ 11. Fungee Jumping:
Students began working on the Jupyter Notebooks for the ED problem in week 10, but our analysis focused solely on the final week’s PDF report. We relied on these reports rather than group discussions, as most teams prioritized solutions over revisiting their problem statements in their discussions (see Table 4). This analysis examines the final product of problem framing, not its evolution over time. For coding results, refer to Appendix A. Figures 5 and 6 present the problem statements from two teams.
@@
Figure 5: Example problem statement from Team-05. The team identified stakeholders (engineering firm, companies), provided the reasoning behind their problem choice, outlined criteria (range and target location) and constraints (specific angle of elevation), modeled the spring to obey Hooke’s law, provided metrics (numerical factors relevant to their problem), stated assumptions (negligible air resistance), and even designated a name for their company. See Appendix A.
@@
Figure 6: Example problem statement from Team-09. The team identified the stakeholders (the company, users (children), parents), provided a reasoning behind their problem choice, stated criteria (material choice, safety) and constraints (safety), did not explicitly mention any numerical factors, and even designated a name for their company. See Appendix A.
@@
An interesting trend emerged: about half the teams assigned creative branding or marketing names to their designs, reflecting an entrepreneurial mindset [135] and engagement beyond mere technical problem-solving. Examples include Fungee Jumping (Team-11’s bungee jumping design), Hookean Bow (Team-05’s bow and arrow), and Duncan’s Toy Chest (Team-09’s bean bag toy). These names showcased students’ creativity and marketing instincts, highlighting an aspect of design thinking that, while valuable, may not be typically prioritized in a physics classroom. See Table 6.
@@
This example also serves to illustrate the variety of ways students chose to frame their problem statement.
About half of the teams explicitly referenced assumptions and approximations in their problem statements. This may reflect an awareness of design complexities or an implicit plan to address these aspects later. A challenge we encountered was distinguishing between assumptions and approximations, as students often used these terms interchangeably. We adopted a flexible approach to accommodate the different ways in which teams chose to communicate. For example, Team-05 stated:
@@
In engineering design-based activities, the role of iterative thinking may need to be emphasized. Following a systematic approach to brainstorming [136] which we had outlined in our recent paper [9], may aid in guiding students to frame their problem statement. Valkenburg and Dorst [137, p.254] emphasize that reflective practices can help rethink and reframe problems, leading to better solutions.
It is also important for students to select problems that not only align with their personal interests but are reasonably simple so that meaningful progress is possible given limited time and resources. Actively seeking input from TAs—can enhance problem framing and engagement.
@@ 5.2 Design-Based Thinking (DBT)
While coding, we considered whether the team intended to refer to safety as a criterion rather than a constraint. One perspective is that safety is a constraint because it represents a non-negotiable requirement that the design must meet. We had a productive discussion on this issue during coding and it is our view that educators take a flexible stance to accommodate students’ styles of thinking and communication. In our experience, this is not an isolated instance of the blurry difference between what may be called a criterion or a constraint. Despite this challenge, these are useful notions to adopt in a design process.
@@
Almost all teams cited sources or referenced online materials (see Figure 19 in Appendix B). A few teams mentioned gaining insights through discussions with their TA, while one team drew inspiration from classroom lectures. Additionally, some teams referenced national regulatory authorities, such as the National Institute of Standards and Technology (Team-08) and the National Highway Traffic Safety Administration (Team-02). While it may be argued that referencing is typically associated with academic rigor rather than design thinking, the practice of referencing plays a complementary role in guiding students towards situating their work, justifying design choices, and informing iterative improvements. Although we did not find literature specifically addressing referencing within design thinking, we believe it serves as a valuable tool for refining ideas and linking them to established knowledge, thus promoting iterative design.
@@
Team-09 chose to include the numerical details while presenting their solutions, while Team-12 found it convenient to employ a python coding cell for this purpose (see Figure 7).
@@
We observed variations in how teams approached iterations. While some engaged in cursory or token iterations, a smaller fraction pursued meaningful refinements. Given that Week 14 was the final week of the semester, some teams may have experienced end-of-semester fatigue [138]. Moreover, more iterations did not always lead to better solutions, highlighting the need for a more systematic approach to guiding students in the iteration process.
@@
To enhance students’ design-based thinking in future iterations, we propose a more detailed rubric, though we acknowledge the risk of it becoming a checklist for students to simply tick off.